OpenStudy (anonymous):

plzzzz help :) ill fan Which of the following represent an expression that has been simplified? a) 2xy + 3xy + x b) 9x - 5 c) 2(x + 1) d) 6x +2 - 17

OpenStudy (rocketisawesome):

I believe it is D?

OpenStudy (rocketisawesome):

But I am not totally sure

OpenStudy (anonymous):

ok can you help me with 2 more plzzz :)

OpenStudy (anthonygarrison875):

i think ist b

OpenStudy (rocketisawesome):

sure

OpenStudy (anonymous):

Which of the following are non-linear expressions? Select all that apply. (There may be more than one correct answer.) Question 2 options: a) x2+x2 b) 13 c) 2+2x−1 d) 3x+5 e) 15x−3+2 f) 3x3−2x2+x

OpenStudy (rocketisawesome):

B

OpenStudy (anthonygarrison875):

it is because all the other ones can be simplified but the second one would be irrational number

OpenStudy (rocketisawesome):

B,A,F I believe

OpenStudy (anonymous):

Which equation is a true statement when x = 2? Question 3 options: a) 4x+2=4x−2 b) 5x+3x−10=4 c) 8(x−1)=8x−8 d) 5+x=−x+2 one more :)

OpenStudy (andrewherpeche):

its b or d